# Simultaneous Khintchine theorem on manifolds in positive characteristics: convergence case

### v1

@@ 4.2. Analysis on Generic Part
$\sigma_{k}:=\begin{pmatrix}0&0&\ldots&0&1\\ 0&0&\ldots&1&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&1&\ldots&0&0\\ 1&0&\ldots&0&0\end{pmatrix}$.
@@
$g^{\star}:=\sigma_{n+1}^{-1}\left(g^{t}\right)^{-1}\sigma_{n+1}\quad\text{for }g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell}).$
@@
$\begin{split}\mathfrak{M}_{0}(s,t):=\left\{\mathbf{x}\in U:\lambda_{1}(g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x})\mathcal{R}^{n+1})<q^{-\frac{(d+2)t-2ns}{2(n+1)}}\right\}\\ =\left\{\mathbf{x}\in U:\lambda_{1}(g_{s,t}^{*}u_{1}^{*}(\mathbf{x})\mathcal{R}^{n+1})<q^{-\frac{(d+2)t-2ns}{2(n+1)}}\right\},\end{split}$ (4.1)
@@
$\mathfrak{M}(s,t)=\bigcup_{\mathbf{x}\in\mathfrak{M}_{0}(s,t)}B(\mathbf{x},q^{-\frac{s+t}{2}}).$
@@
We first estimate the number of points in the generic part of the manifold. Towards this end, we prove the following proposition.
Proposition 2.
Let $\mathbf{f}$ and $U$ be as in Proposition 1. Then, for any $s\in\mathbb{N}$, for any ball $B\subset U$ , and all sufficiently large $t$, we have
$\mathcal{N}_{\bm{\Theta}}(B\setminus\mathfrak{M}(s,t);s,t)\ll q^{(d+1)t}q^{-ms}\mathcal{L}_{d}(B).$ (4.2)
The proof of Proposition 2 relies on the following counting estimate:
@@
Let $\mathbf{f}$ and $U$ be as in Proposition 1. Also let $B$ be an open ball inside $U\subset\mathcal{K}^{d}$. Then, for all $\mathbf{x}_{0}\in(U\setminus\mathfrak{M}(s,t))\cap B$ and $t>0$ sufficiently large,
$\mathcal{N}_{\bm{\Theta}}(B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})\cap B;s,t)\ll q^{\frac{d}{2}(t+s)+(t-ns)}.$
@@
Without loss of generality, assume that $\mathcal{N}_{\bm{\Theta}}(B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})\cap B;s,t)>1$. Then for $i=1,2$ there exist $(\mathbf{P}_{i},Q_{i})\in\mathcal{R}^{n+1}$ and $\mathbf{x}_{i}\in B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})$, such that
$\left\|\mathbf{f}(\mathbf{x}_{i})-\frac{\mathbf{P}_{i}+\bm{\Theta}}{Q_{i}}\right\|<q^{-(s+t)}.$ (4.3)
Hence, by Lemma 4.2, we have for $h=1,2$, and for every $i=1,\dots,d,$, we have
$|Q_{h}x_{0,i}-P_{h,i}-\Theta_{i}|<\max\{q^{-s},q^{t-\frac{s+t}{2}}\}.$ (4.4)
@@
$\begin{split}\left|Q_{h}f_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Q_{h}x_{0,i}-P_{h,i}-\Theta_{i})-P_{h,d+j}-\Theta_{d+j}\right|\leq\max\{1,M\}q^{-s}.\end{split}$ (4.5)
Let $Q=Q_{2}-Q_{1}$ and let $\mathbf{P}=\mathbf{P}_{2}-\mathbf{P}_{1}$, so that $|Q|\leq q^{t}$. Hence, by subtracting (4.4) from each other with $h=1,2$, we have
$|Qx_{0,i}-P_{i}|<\max\left\{q^{-s},q^{t-\frac{s+t}{2}}\right\}=\max\left\{q^{-s},q^{\frac{t-s}{2}}\right\}=q^{\frac{t-s}{2}},$ (4.6)
whenever $t$ is large enough. Similarly, by subtracting (4.5) from each other with $h=1,2$, we have
$\begin{split}\left|Qf_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Qx_{0,i}-P_{i})-P_{d+j}\right|\leq\max\{1,M\}q^{-s}.\end{split}$ (4.7)
@@
$u_{1}(\mathbf{x}_{0})\begin{pmatrix}-\mathbf{P}\sigma_{n}\\ Q\end{pmatrix}\in\left(\left[\max\{1,M\}q^{-s}\right]^{m}\times\left[q^{\frac{t-s}{2}}\right]^{d}\times[q^{t}]\right)\cap u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1}.$
@@
$g_{s,t}u_{1}(\mathbf{x}_{0})\begin{pmatrix}-\mathbf{P}\sigma_{n}\\ Q\end{pmatrix}\in\overbrace{\left(\max\{1,M\}\left(\left[q^{\frac{(d+2)t-2ns}{2(n+1)}}\right]^{n+1}\right)\right)}^{\Omega}\cap g_{s,t}u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1}.$
Hence, one has $\lambda_{1}(g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x}_{0})\mathcal{R}^{n+1})=\lambda_{1}(g_{s,t}^{*}u_{1}^{*}(\mathbf{x}_{0})\mathcal{R}^{n+1})\geq q^{-\frac{(d+2)t-2ns}{2(n+1)}}.$ Now, it follows from duality (Theorem 3.3) that $\lambda_{n+1}(g_{s,t}u_{1}(\mathbf{x_{0}}))\mathcal{R}^{n+1})\ll_{n}q^{\frac{(d+2)t-2ns}{2(n+1)}}$. Therefore, by applying Lemma 3.6 with $c=\frac{(d+2)t-2ns}{2(n+1)}$, for every sufficiently large $t>0$, we have
$\#g_{s,t}u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1}\cap\max\{1,M\}\left[q^{\frac{(d+2)t-2ns}{2(n+1)}}\right]^{n+1}$ $\ll\frac{q^{\frac{(d+2)t-2ns}{2}}}{q^{\frac{(d+2)t+2s}{2}-\frac{d+2}{2}(s+t)}}$
$=q^{\left(\frac{d}{2}+1\right)t-s\left(n-\frac{d}{2}\right)}$
$=q^{\frac{d}{2}(t+s)+(t-ns)}.$
The following lemma, coupled with Lemma 4.3, now yields Proposition 2:
@@
For all sufficiently large $t>0$, we have
$\mathcal{N}_{\mathbf{\Theta}}(B\setminus\mathfrak{M}(s,t);s,t)\leq q^{\frac{d(t+s)}{2}}\mathcal{L}_{d}(B)\max_{\mathbf{x}_{0}\in B\setminus\mathfrak{M}(s,t)}\mathcal{N}_{\mathbf{\Theta}}(B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})\cap B;s,t).$
@@
The proof of this Lemma is similar to that of [BY23, Lemma 5.4]. Firstly, one can cover $B$ with at most $q^{\frac{d}{2}(s+t)}\mathcal{L}_{d}(B)$ balls of radius $q^{-\frac{s+t}{2}}$. If one of these balls intersects $\mathfrak{M}(s,t)$, then there exists $\mathbf{x}_{0}\in\mathfrak{M}(s,t)$ such that the ball coincides with $B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})$. Therefore,
$\mathcal{N}_{\mathbf{\Theta}}(B\setminus\mathfrak{M}(s,t);s,t)\leq q^{\frac{d}{2}(t+s)}\mathcal{L}_{d}(B)\max_{\mathbf{x}_{0}\in B\setminus\mathfrak{M}(s,t)}\mathcal{N}_{\mathbf{\Theta}}(B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})\cap B;s,t).$
4.3. Analysis on Special Part
To deal with the special part of the manifold, for $r,s^{\prime},t_{1},\dots,t_{n}\in\mathbb{Z}$ and for any ball $B\subset\mathcal{K}^{d}$, define
$\mathfrak{G}_{\mathbf{f}}(r,s^{\prime},t_{1},\dots,t_{n})=\left\{\mathbf{x}\in B:\exists\,\,\mathbf{P}=(P_{1},\dots,P_{n})\in\mathcal{R}^{n},Q_{0}\in\mathcal{R}\left|\begin{array}[]{lcl}|\mathbf{f}(\mathbf{x})\cdot\mathbf{P}+Q_{0}|<q^{-r}\\[8.61108pt] \|\nabla(\mathbf{f}(\mathbf{x})\cdot\mathbf{P})\|<q^{s^{\prime}}\\[8.61108pt] |P_{i}|<q^{t_{i}},\,\,\,i=1,\dots,n\end{array}\right.\right\}$.
We use the following theorem by Das and Ganguly [DG22], which bounds the measure of the set $\mathfrak{G}_{\mathbf{f}}(r,s^{\prime},t_{1},\dots,t_{n})$.
Theorem 4.5 ([DG22, Theorem 6.1]).
Suppose $U$ is an open subset of $\mathcal{K}^{d}$ and $\mathbf{f}:U\rightarrow\mathcal{K}^{n}$ satisfies (II), (III) and (IV). Then for any $\mathbf{x}_{0}\in U,$ one can find a neighbourhood $V\subseteq U$ of $\mathbf{x}_{0}$ and $\alpha>0$ with the following property: for any ball $B\subseteq V,$ there exists $E>0$ such that for any choice of $r,s^{\prime},t_{1},\dots,t_{n}\in\mathbb{Z}$ with $r\geq 0,$ $t_{1},\dots,t_{n}\geq 1,$ and $s^{\prime}+\sum_{i}t_{i}-r-\max_{i}t_{i}<0$ one has
$\mathcal{L}_{d}\left(\mathfrak{G}_{\mathbf{f}}(r,s^{\prime},t_{1},\dots,t_{n})\right)\leq E\gamma^{\alpha}\mathcal{L}_{d}(B),$ (4.8)
where $\gamma:=\max\left(q^{-r},q^{\frac{s^{\prime}+\sum_{i}t_{i}-r-\max_{i}t_{i}}{n+1}}\right)$.
We now prove that the special part of the manifold $\mathfrak{M}(s,t)$ is contained inside the set $\mathfrak{G}_{f}(\widetilde{C}t,\widetilde{C}\frac{s-t}{2},\widetilde{C}s,\dots,\widetilde{C}s)$ for sufficiently large $t$ and some constant $\widetilde{C}$.
@@ 4.3. Analysis on Special Part
Let $\mathbf{f}$ and $U$ be as in Proposition 1. For $t$ large enough, there exists a constant $\widetilde{C}\in\mathbb{N}$ (depending on $M,d$ and $n$) such that we have
$\begin{split}\mathfrak{M}(s,t)\subseteq\mathfrak{G}_{\mathbf{f}}\left(\widetilde{C}t,\widetilde{C}\frac{s-t}{2},\widetilde{C}s,\dots,\widetilde{C}s\right).\end{split}$
@@
$g_{s,t}^{\star}=\operatorname{diag}\left(T^{s+t-\frac{(d+2)t+2s}{2(n+1)}},T^{\frac{s+t}{2}-\frac{(d+2)t+2s}{2(n+1)}},\dots,T^{\frac{s+t}{2}-\frac{(d+2)t+2s}{2(n+1)}},T^{-\frac{(d+2)t+2s}{2(n+1)}},\dots,T^{-\frac{(d+2)t+2s}{2(n+1)}}\right)$
and $u_{1}^{\star}(\mathbf{x}_{0})=\begin{pmatrix}1&-\mathbf{x}_{0}&-f(\mathbf{x}_{0})\\ &\mathrm{I}_{d}&J(\mathbf{x}_{0})\\ &&\mathrm{I}_{m}\end{pmatrix}$. Then, by (4.9), there exists some $(Q_{0},\mathbf{P}):=(Q_{0},\mathbf{P}^{(d)},\mathbf{P}^{(m)})=(Q_{0},P_{1}^{(d)},\dots,P_{d}^{(d)},P_{1}^{(m)},\dots,P_{m}^{(m)})\in\mathcal{R}^{n+1}\setminus\{0\}$, such that
$\begin{array}[]{lcl}|Q_{0}+\mathbf{P}\cdot\mathbf{f}(\mathbf{x}_{0})|<q^{-t},\\[8.61108pt] \left|{P}^{(d)}_{i}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}_{0}){P}^{(m)}_{j}\right|<q^{\frac{s-t}{2}},\,\text{ for }i=1,\dots,d\\[8.61108pt] |{P}^{(m)}_{j}|<q^{s},\,\text{ for }j=1,\dots,m.\end{array}$ (4.10)
@@
$\begin{split}\left|{P}^{(d)}_{i}\right|&\leq\max\left\{\left|{P}^{(d)}_{i}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}_{0}){P}^{(m)}_{j}\right|,\max_{j=1,\dots,m}\left|\partial_{i}f_{j}(\mathbf{x}_{0}){P}^{(m)}_{j}\right|\right\}\\ &\leq\max\{q^{-t},Mq^{s}\}\\ &\leq\max\{1,M\}q^{s}.\end{split}$ (4.11)
If $\mathbf{x}\in\mathfrak{M}(s,t)$, then there exists $\mathbf{x}_{0}\in\mathfrak{M}_{0}(s,t)$, such that $\|\mathbf{x}-\mathbf{x}_{0}\|<q^{-\frac{s+t}{2}}$. By Taylor’s expansion (Lemma 2.2) of $\mathbf{f}$ around $\mathbf{x}_{0}$, the boundedness of second-order partial difference quotients of $f_{i}$ (i.e., (IV)), the ultrametric inequality (4.10), and (4.11), we obtain
$|Q_{0}+\mathbf{P}\cdot\mathbf{f}(\mathbf{x})|\leq\max\left\{q^{-t},q^{-\frac{s+t}{2}}q^{\frac{s-t}{2}},Mq^{s}\,\left(q^{-(\frac{s+t}{2})}\right)^{2}\right\}<\max\{1,M\}q^{-t}.$ (4.12)
Also, using the inequality (4.11) and Taylor’s expansion (Lemma 2.2) of $\partial_{i}f_{j}$ around the point $\mathbf{x}_{0}$, we obtain
$\left|\mathbf{P}_{i}^{(d)}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x})\mathbf{P}_{j}^{(m)}\right|$ $\leq$ $\max\left\{\left|\mathbf{P}_{i}^{(d)}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}_{0})\mathbf{P}_{j}^{(m)}\right|,\max_{j=1,\dots,m}\left|\partial_{i}f_{j}(\mathbf{x})-\partial_{i}f_{j}(\mathbf{x}_{0})\right|\cdot\left|\mathbf{P}_{j}^{(m)}\right|\right\}$ (4.13)
$\leq$ $\max\left\{q^{\frac{s-t}{2}},Mq^{-\frac{s+t}{2}}q^{s}\right\}$
$\leq$ $\max\{1,M\}q^{\frac{s-t}{2}}.$
Finally, combining (4.12), (4.13), (4.10), and (4.11), there exists a constant $\widetilde{C}\in\mathbb{N}$ that depends only on $M$, $d$, and $n$, such that
$\begin{split}\mathfrak{M}(s,t)\subseteq\mathfrak{G}_{\mathbf{f}}\left(\widetilde{C}t,\widetilde{C}\frac{s-t}{2},\widetilde{C}s,\dots,\widetilde{C}s\right).\end{split}$
@@
Let $U\subseteq\mathcal{K}^{d}$ be an open set and $\mathbf{f}:U\rightarrow\mathcal{K}^{n}$ be of the form $\mathbf{f}(\mathbf{x})=(\mathbf{x},\bm{f}(\mathbf{x}))$ such that it satisfies the hypothesis of Proposition 1. Then there exists $C,\alpha>0$ such that
$\mathcal{L}_{d}(\mathfrak{M}(s,t)\cap B)\leq C\max\left\{q^{-t},q^{\frac{(2n-1)s-3t}{2(n+1)}}\right\}^{\alpha}\mathcal{L}_{d}(B),$
whenever $t$ is large enough.

### v2

@@ 4.2. Analysis on the Generic Part
$\sigma_{k}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\begin{pmatrix}0&0&\ldots&0&1\\ 0&0&\ldots&1&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&1&\ldots&0&0\\ 1&0&\ldots&0&0\end{pmatrix}$.
@@
$g^{\star}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\sigma_{n+1}^{-1}\left(g^{t}\right)^{-1}\sigma_{n+1}\quad\text{for }g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell}).$
@@
$\begin{split}\mathfrak{M}_{0}(s,t){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\left\{\mathbf{x}\in U:\lambda_{1}(g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x})\mathcal{R}^{n+1})<q^{-\frac{(d+2)t-2ns}{2(n+1)}}\right\}\\ =\left\{\mathbf{x}\in U:\lambda_{1}(g_{s,t}^{*}u_{1}^{*}(\mathbf{x})\mathcal{R}^{n+1})<q^{-\frac{(d+2)t-2ns}{2(n+1)}}\right\},\end{split}$ (4.1)
@@
$\mathfrak{M}(s,t)=\bigcup_{\mathbf{x}\in\mathfrak{M}_{0}(s,t)}B\left(\mathbf{x},q^{-\frac{s+t}{2}}\right).$
@@
We first estimate the number of points in the manifold’s generic part. Towards this end, we prove the following proposition.
Proposition 4.1.
Let $\mathbf{f}$ and $U$ be as in Proposition 1.1. Then, given any ball $B\subseteqq U$, the following holds for sufficiently large $t\in\mathbb{N}$:
$\mathcal{N}_{\bm{\Theta}}(B\setminus\mathfrak{M}(s,t);s,t)\ll q^{(d+1)t}q^{-ms}\mathcal{L}_{d}(B),\penalty 10000\ \forall s\in\mathbb{N}.$ (4.2)
The proof of Proposition 4.1 relies on the following counting estimate:
@@
Let $\mathbf{f}$ and $U$ be as in Proposition 1.1. Consider $s,t\in\mathbb{N}$, and an open ball $B\subseteqq U\subseteqq\mathcal{K}^{d}$. Then, for all $\mathbf{x}_{0}\in(U\setminus\mathfrak{M}(s,t))\cap B$,
$\mathcal{N}_{\bm{\Theta}}\left(B\left(\mathbf{x}_{0},q^{-\frac{s+t}{2}}\right)\cap B;s,t\right)\ll q^{\frac{d}{2}(t+s)+(t-ns)}.$
@@
Let $\mathbf{x}_{0}=(x_{1,0},\dots,x_{d,0})$. Without loss of generality, assume that $\mathcal{N}_{\bm{\Theta}}\left(B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})\cap B;s,t\right)$ is greater than $1$. Then for $i=1,2$ there exist $(\mathbf{P}_{i},Q_{i})\in\mathcal{R}^{n}\times\mathcal{R}\setminus\{0\}$ and $\mathbf{x}_{i}\in B\left(\mathbf{x}_{0},q^{-\frac{s+t}{2}}\right)$, such that $\left\|\mathbf{f}(\mathbf{x}_{i})-\frac{\mathbf{P}_{i}+\bm{\Theta}}{Q_{i}}\right\|<q^{-(s+t)}$ with $|Q_{i}|=q^{t}.$ Hence, by Lemma 4.2, we have for $h=1,2$, and for every $k=1,\dots,d,$, we have
$|Q_{h}x_{k,0}-P_{h,k}-\Theta_{k}|<\max\left\{q^{-s},q^{t}q^{-\frac{s+t}{2}}\right\}=q^{\frac{t-s}{2}}.$ (4.3)
@@
$\begin{split}\left|Q_{h}f_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Q_{h}x_{i,0}-P_{h,i}-\Theta_{i})-P_{h,d+j}-\Theta_{d+j}\right|\leq Mq^{-s}.\end{split}$ (4.4)
Let $Q=Q_{2}-Q_{1}$ and $\mathbf{P}=\mathbf{P}_{2}-\mathbf{P}_{1}=(P_{1},\dots,P_{n})$ so that $|Q|\leq q^{t}$. Hence, by subtracting (4.3) from each other with $h=1,2$, we have
$|Qx_{i,0}-P_{i}|<q^{\frac{t-s}{2}}\,\,\,\,\text{for }i=1,\dots,d$ (4.5)
Similarly, by subtracting (4.4) from each other with $h=1,2$, we have
$\begin{split}\left|Qf_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Qx_{i,0}-P_{i})-P_{d+j}\right|\leq Mq^{-s}\quad\text{for }j=1,\dots,m.\end{split}$ (4.6)
@@
$u_{1}(\mathbf{x}_{0})\begin{pmatrix}-\mathbf{P}\sigma_{n}\\ Q\end{pmatrix}\in\left(\left[Mq^{-s}\right]^{m}\times\left[q^{\frac{t-s}{2}}\right]^{d}\times[q^{t}]\right)\cap u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1}.$
@@
$g_{s,t}u_{1}(\mathbf{x}_{0})\begin{pmatrix}-\mathbf{P}\sigma_{n}\\ Q\end{pmatrix}\in{\left(M\left[q^{\frac{(d+2)t-2ns}{2(n+1)}}\right]^{n+1}\right)}\cap g_{s,t}u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1}.$
Hence, one has $\lambda_{1}(g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x}_{0})\mathcal{R}^{n+1})=\lambda_{1}(g_{s,t}^{*}u_{1}^{*}(\mathbf{x}_{0})\mathcal{R}^{n+1})\geq q^{-\frac{(d+2)t-2ns}{2(n+1)}}.$ Now, it follows from duality (Theorem 3.3) that $\lambda_{n+1}(g_{s,t}u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1})\ll_{n}q^{\frac{(d+2)t-2ns}{2(n+1)}}$. Therefore, by applying Lemma 3.7 with $c=\frac{(d+2)t-2ns}{2(n+1)}$, we have
$g_{s,t}u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1}\cap M\left[q^{\frac{(d+2)t-2ns}{2(n+1)}}\right]^{n+1}\ll\frac{q^{\frac{(d+2)t-2ns}{2}}}{q^{\frac{(d+2)t+2s}{2}-\frac{d+2}{2}(s+t)}}=q^{\frac{d}{2}(t+s)+(t-ns)}.$
The following lemma, coupled with Lemma 4.3, now yields Proposition 4.1:
@@
For all sufficiently large $t\in\mathbb{N}$, we have the following for any $s\in\mathbb{N}$:
$\mathcal{N}_{\bm{\Theta}}(B\setminus\mathfrak{M}(s,t);s,t)\leq q^{\frac{d(t+s)}{2}}\mathcal{L}_{d}(B)\max_{\mathbf{x}_{0}\in B\setminus\mathfrak{M}(s,t)}\mathcal{N}_{\bm{\Theta}}\left(B\left(\mathbf{x}_{0},q^{-\frac{s+t}{2}}\right)\cap B;s,t\right).$
@@
Choose $t\in\mathbb{N}$ large enough so that $q^{-\frac{t}{2}}$ is at most the radius of $B$. Then any open ball centred at a point in $B$ and having radius not exceeding $q^{-\frac{t}{2}}$ must be contained in $B$. From this, it follows that, when $t\gg 1$, for all $s\in\mathbb{N}$, $B$ can be written as a disjoint union of open balls having radius $q^{-\frac{s+t}{2}}$, and clearly the number of such balls in the union is at most $q^{\frac{d}{2}(s+t)}\mathcal{L}_{d}(B)$. If any of these balls intersects $\mathfrak{M}(s,t)$, then that ball coincides with $B\left(\mathbf{x}_{0},q^{-\frac{s+t}{2}}\right)$ for some $\mathbf{x}_{0}\in\mathfrak{M}(s,t)$. Therefore,
$\mathcal{N}_{\bm{\Theta}}(B\setminus\mathfrak{M}(s,t);s,t)\leq q^{\frac{d}{2}(t+s)}\mathcal{L}_{d}(B)\max_{\mathbf{x}_{0}\in B\setminus\mathfrak{M}(s,t)}\mathcal{N}_{\bm{\Theta}}\left(B\left(\mathbf{x}_{0},q^{-\frac{s+t}{2}}\right)\cap B;s,t\right).$
4.3. Analysis on the Special Part
To deal with the special part of the manifold, for $r,s^{\prime},t_{1},\dots,t_{n}\in\mathbb{Z}$ and for any ball $B\subseteqq\mathcal{K}^{d}$, define
$\mathfrak{G}_{\mathbf{f}}(r,s^{\prime},t_{1},\dots,t_{n})=\left\{\mathbf{x}\in B:\exists\,\,\mathbf{P}=(P_{1},\dots,P_{n})\in\mathcal{R}^{n},Q\in\mathcal{R}\left|\begin{array}[]{lcl}|\mathbf{f}(\mathbf{x})\cdot\mathbf{P}+Q|<q^{-r}\\[8.61108pt] \|\nabla(\mathbf{f}(\mathbf{x})\cdot\mathbf{P})\|<q^{s^{\prime}}\\[8.61108pt] |P_{i}|<q^{t_{i}},\,\,\,i=1,\dots,n\end{array}\right.\right\}$.
We use the following theorem by Das and Ganguly [DG25], which bounds the measure of the set $\mathfrak{G}_{\mathbf{f}}(r,s^{\prime},t_{1},\dots,t_{n})$.
Theorem 4.5 ([DG25, Theorem 6.1]).
Suppose $U$ is an open subset of $\mathcal{K}^{d}$ and $\mathbf{f}:U\rightarrow\mathcal{K}^{n}$ which satisfies ($\mathfrak{A}$.1)-($\mathfrak{A}$.4). Then for any $\mathbf{x}_{0}\in U,$ one can find a neighbourhood $V\subseteqq U$ of $\mathbf{x}_{0}$ and $\alpha>0$ with the following property: for any ball $B\subseteqq V,$ there exists $E>0$ such that for any choice of $r,s^{\prime},t_{1},\dots,t_{n}\in\mathbb{Z}$ with $r\geq 0,$ $t_{1},\dots,t_{n}\geq 1,$ and $s^{\prime}+\sum_{i}t_{i}-r-\max_{i}t_{i}<0$ one has
$\mathcal{L}_{d}\left(\mathfrak{G}_{\mathbf{f}}(r,s^{\prime},t_{1},\dots,t_{n})\right)\leq E\gamma^{\alpha}\mathcal{L}_{d}(B),\text{ where }\gamma{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\max\left(q^{-r},q^{\frac{s^{\prime}+\sum_{i}t_{i}-r-\max_{i}t_{i}}{n+1}}\right).$
@@ 4.3. Analysis on the Special Part
Let $\mathbf{f}$ and $U$ be as in Proposition 1.1. Then there exists a constant $\widetilde{C}\in\mathbb{N}$, possibly depending on $M,d$ and $n$, such that for all $t\in\mathbb{N}$, we have
$\begin{split}\mathfrak{M}(s,t)\subseteqq\mathfrak{G}_{\mathbf{f}}\left(\widetilde{C}t,\widetilde{C}\lceil(s-t)/2\rceil,\widetilde{C}s,\dots,\widetilde{C}s\right).\end{split}$
@@
$g_{s,t}^{\star}=\operatorname{diag}\left(T^{s+t-\frac{(d+2)t+2s}{2(n+1)}},\underbrace{T^{\frac{s+t}{2}-\frac{(d+2)t+2s}{2(n+1)}},\dots,T^{\frac{s+t}{2}-\frac{(d+2)t+2s}{2(n+1)}}}_{d},\overbrace{T^{-\frac{(d+2)t+2s}{2(n+1)}},\dots,T^{-\frac{(d+2)t+2s}{2(n+1)}}}^{m}\right)$
and $u_{1}^{\star}(\mathbf{x}_{0})=\begin{pmatrix}1&-\mathbf{x}_{0}&-\bm{f}(\mathbf{x}_{0})\\ &\mathrm{I}_{d}&J(\mathbf{x}_{0})\\ &&\mathrm{I}_{m}\end{pmatrix}$. Then, by (4.7), there exists some $(Q_{0},\mathbf{P}){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}(Q_{0},\mathbf{P}^{(d)},\mathbf{P}^{(m)})=\left(Q_{0},P_{1}^{(d)},\dots,P_{d}^{(d)},P_{1}^{(m)},\dots,P_{m}^{(m)}\right)\in\mathcal{R}^{n+1}\setminus\{0\}$, such that
$\begin{array}[]{lcl}|Q_{0}+\mathbf{P}\cdot\mathbf{f}(\mathbf{x}_{0})|<q^{-t},\\[8.61108pt] \left|{P}^{(d)}_{i}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}_{0}){P}^{(m)}_{j}\right|<q^{\frac{s-t}{2}},\,\text{ for }i=1,\dots,d,\\[8.61108pt] |{P}^{(m)}_{j}|<q^{s},\,\text{ for }j=1,\dots,m.\end{array}$ (4.8)
@@
$\left|{P}^{(d)}_{i}\right|\leq\max\left\{\left|{P}^{(d)}_{i}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}_{0}){P}^{(m)}_{j}\right|,\max_{j=1,\dots,m}\left|\partial_{i}f_{j}(\mathbf{x}_{0}){P}^{(m)}_{j}\right|\right\}\leq Mq^{s}.$ (4.9)
If $\mathbf{x}\in\mathfrak{M}(s,t)$, then there exists $\mathbf{x}_{0}\in\mathfrak{M}_{0}(s,t)$, such that $\|\mathbf{x}-\mathbf{x}_{0}\|<q^{-\frac{s+t}{2}}$. Considering the Taylor expansion about $\mathbf{x}_{0}$, it follows from Lemma 2.2, ($\mathfrak{A}$.4), (4.8), and (4.9) that
$|Q_{0}+\mathbf{P}\cdot\mathbf{f}(\mathbf{x})|\leq\max\left\{q^{-t},Mq^{-\frac{s+t}{2}}q^{\frac{s-t}{2}},Mq^{s}\,\left(q^{-(\frac{s+t}{2})}\right)^{2}\right\}<\max\{1,M\}q^{-t}.$ (4.10)
Again using (4.9) and the first-order Taylor expansion of $\partial_{i}f_{j}$ about $\mathbf{x}_{0}$ (4.9), we get
$\left|{P}_{i}^{(d)}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}){P}_{j}^{(m)}\right|$ $\leq\max\left\{\left|{P}_{i}^{(d)}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}_{0}){P}_{j}^{(m)}\right|,\max_{j=1,\dots,m}\left|\partial_{i}f_{j}(\mathbf{x})-\partial_{i}f_{j}(\mathbf{x}_{0})\right|\cdot\left|{P}_{j}^{(m)}\right|\right\}$
$\leq\max\left\{q^{\frac{s-t}{2}},Mq^{-\frac{s+t}{2}}q^{s}\right\}\leq\max\{1,M\}q^{\frac{s-t}{2}}.$ (4.11)
Finally, combining (4.10), (4.3), (4.8), and (4.9), there exists a constant $\widetilde{C}\in\mathbb{N}$ that depends only on $M$, $d$, and $n$, such that the set $\mathfrak{G}_{\mathbf{f}}\left(\widetilde{C}t,\widetilde{C}\lceil(s-t)/2\rceil,\widetilde{C}s,\dots,\widetilde{C}s\right)$ contains $\mathfrak{M}(s,t)$.
@@
Let $U\subseteqq\mathcal{K}^{d}$ be an open set and $\mathbf{f}:U\rightarrow\mathcal{K}^{n}$ be of the form $\mathbf{f}(\mathbf{x})=(\mathbf{x},\bm{f}(\mathbf{x}))$ such that it satisfies the hypothesis of Proposition 1.1. Then for every $\mathbf{x}_{0}\in U$ there exists a ball $B$ centered at $\mathbf{x}_{0}$ and constants $C,\alpha>0$ such that
$\mathcal{L}_{d}(\mathfrak{M}(s,t)\cap B)\leq Cq^{\frac{\alpha}{n+1}\left(ns-\frac{3t}{2}\right)}\mathcal{L}_{d}(B),\text{ whenever $t\in\mathbb{N}$ is large enough and }s\leq\frac{\kappa t}{n}.$